GE’s stock had a required return of 15% last year, when the risk-free rate was 4% and the market risk premium was 6%. Then an increase in investor risk aversion caused the market risk premium to rise by 3%. The risk-free rate and the firm's beta remain unchanged. What is the company's new required rate of return?

24.18%

22.15%

20.50%

18.23%

15.87%
Answer #1

Required rate=risk free rate+Beta*market risk premium

15=4+Beta*6

Beta=(15-4)/6

=1.833(Approx).

New market risk premium=(6+3)=9%

Hence new required rate=4+(1.833*9)

=20.5%